40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 11063     Entry time: Wed Feb 25 04:21:58 2015
Author: ericq 
Type: Update 
Category: CDS 
Subject: Some model updates 

I changed the suspension library block to acquire the SUS_[optic]_LSC_OUT channels at 16k for sensing matrix investigations. We could save the FB some load by disabling these and oplev channels in the mode cleaner optic suspensions. 

I removed nonexistant PDs from c1cal, to try and speed it up from its constantly overflowing state. It's still pretty bad, but under 60us most of the time. 

I also cleaned out the unused IPCs for simulated plant stuff from c1scx and c1sus, to get rid of red blinkeys. 

ELOG V3.1.3-